and repeats the measurement 3245 times and records the mean of his 3245 measurements What is the standard deviation (20.01) of suans mean resut? (That is, i uan kapt on Making 3245 measurements and averaging them, what would be the standard deviation of all his z s?) so that the standard deviation of F is 37 Round your answer to the nearest whole number.)
Answer #1

Answer

Here by the problem,

Juan makes a measurement in a chemistry laboratory and record the results in his lab report. The standard deviation of the students's lab measurements is s=13 miligrams. Juan repeats the measurement3245 times and records of his 3245 measurements.

Then the standard deviation of Juan's mean result be,

SD(\bar{x})=\frac{s}{\sqrt{n}}=\frac{13}{\sqrt{3245}}=0.228\approx 0.23

Now in order to getthe standard deviation of \bar{x} as 3, if here we have m samples then,

SD(\bar{x})=\frac{s}{\sqrt{m}}=3

\Rightarrow m=\frac{s^2}{9}=\frac{13^2}{9}=18.78\approx 19

Hence the number of required sample size.

Hence the answer............

Thankyou.............